# Attention Guidance Mechanism for Handwritten Mathematical Expression Recognition

### 04 Mar 2024 v1

@@ Abstract
Handwritten mathematical expression recognition (HMER) is challenging in OCR tasks due to the complex layouts of mathematical expressions, suffering from issues including over-parsing and under-parsing. To solve these, previous methods utilize historical attention weights to improve the attention mechanism. However, this approach has limitations in addressing under-parsing since it cannot correct the erroneous attention on image regions that should be parsed at subsequent decoding steps. When this happens, the attention module incorporates future context into the current decoding step, thus confusing the alignment process. To address this issue, we propose an attention guidance mechanism to explicitly suppress attention weights in irrelevant regions and enhance ones in appropriate regions, thereby inhibiting access to information outside the intended context. Depending on the type of attention guidance, we devise two complementary approaches to refine attention weights: self-guidance that coordinates attention of multiple heads and neighbor-guidance that integrates attention from adjacent time steps. Experiments show that our method outperforms existing state-of-the-art methods, achieving expression recognition rates of 60.75% / 61.81% / 63.30% on the CROHME 2014 / 2016 / 2019 datasets.
@@ 1 Introduction
Handwritten mathematical expression recognition (HMER) aims to translate mathematical expression images into corresponding  sequences. HMER is a fundamental OCR task in educational scenarios, which has led to many real-world applications, including automatic scoring and online education. However, HMER is challenging because mathematical expressions are laid out in two dimensions, which is different from natural language texts including English, Chinese, etc. An HMER algorithm is expected to analyze the structure of an expression while recognizing the symbols (1). Over the past few years, HMER methods have benefited from deep neural networks and have made significant progress (2; 3; 4; 5; 6; 7; 8; 9; 10; 11; 12). Most prevalent methods treat HMER as a sequence-to-sequence task and follow the attention-based encoder-decoder framework. In this framework, the encoder first extracts the visual features of the input image as a sequence, and then the decoder generates the target sequence by attending to the encoded feature sequence (2; 4; 5) at each decoding step. The attention mechanism attends to image regions at each decoding step based on the hidden state of the last step.
@@
Beyond the hidden state, many HMER methods (2; 6; 3; 8; 9) exploit historical attention weights to prompt the attention module to focus on unattended image areas, called coverage mechanism. The coverage mechanism is proposed to solve the over-parsing and under-parsing problems. Over-parsing occurs when a particular symbol region is parsed multiple times unnecessarily, whereas under-parsing happens when a symbol region is overlooked and not parsed at all (2). The coverage mechanism successfully alleviates over-parsing but has limitations in addressing under-parsing. Fig. 1 illustrates a series of attended image regions from a state-of-the-art coverage-based system (6). As we can see, the superscripts of “$j^{2}$” and “$q^{2}$” are activated at the same time. As a result, features of another “2” that should be decoded in the future are incorporated in the current context, causing “$q^{2}$” to be omitted in the decoding process.
@@
We refer to the above observation context leakage phenomenon: if the attention module attends to image regions that should be parsed after the target ones, context from the future is incorporated into the attention result, causing symbols between the target region and the incorrectly attended region to tend to be omitted. The coverage mechanism cannot solve this phenomenon since it is only aware of past alignment information. Different from the multi-activation phenomenon in (13; 14), the context leakage phenomenon emphasizes the incorrect predictions caused by the activations of image regions should be parsed after the target ones. To address this issue, we propose an attention guidance mechanism which aims to explicitly suppress attention weights in irrelevant regions and enhance ones in appropriate regions. Specifically, this mechanism includes two complementary approaches: self-guidance and neighbor-guidance.
@@
Self-guidance aims to refine correlation maps by eliminating inconsistencies of different attention heads. In multi-head attention, each head learns its own set of linear projection matrices for queries, keys, and values. This independence allows the model to capture a richer variety of information from different perspectives. Although different heads attend to image areas in different subspaces, the attended symbol in the input image should be consistent. However, due to the presence of symbols with similar visual features, multiple symbols may be attended by different heads. The incorrectly attended symbols are the cause of the context leakage phenomenon. Therefore, inconsistent attention needs to be refined. We achieve this through a guidance map indicating inconsistencies among attention heads and a linear projection to adjust the guidance weights.
While self-guidance is performed in each single decoding step, attention from adjacent time steps can also be involved in the attention guidance mechanism. For HMER methods with a stacked decoder (15; 5; 6; 10), the query of the attention module in each decoder layer is determined by the output of its previous layer. Fig. 2 illustrates how a stacked decoder generates the final attention regions. As we can see, only the output layer focuses on the symbol to be generated. The middle layers first calculate attention weights based on the previous symbol instead of directly aligning with the current one. In other words, each symbol is aligned by its previously decoded neighbor. Different HMER methods (5; 6) with different decoder layers ($L\geq 2$) follow this process consistently. Moreover, a similar phenomenon has also been found in image captioning systems (16; 17).
@@
Based on the above observation, we propose another attention guidance approach, neighbor-guidance. Neighbor-guidance aims to leverage the final attention weights of the previous decoding step to enhance appropriate attention for middle layers at the current step. For example, in Fig. 2, attention weights for generating “$x$” in layer $L$ can be reused to refine the attention weights of layer $L-2$ at decoding step 2, where the three “$x$” are attended simultaneously and the context leakage phenomenon occurs. With the help of neighbor-guidance, appropriate attention weights of middle layers are enhanced, thus improving the attention quality of both the current and final layers. Neighbor-guidance is implemented similarly to the self-guidance module but changes the attention guidance from self-generated attention to the attention of the last decoding step and can be readily employed without training. Moreover, different from self-guidance, neighbor-guidance can be applied only to middle layers and facilitates information propagation between adjacent time steps.
@@
    We propose the phenomenon of context leakage, which cannot be solved by the widely used coverage mechanism. To solve this issue, we propose a general attention guidance mechanism that aims to suppress attention weights in irrelevant regions and enhance ones in appropriate regions explicitly.
@@
    We devise self-guidance to refine the raw attention by seeking consensus among different attention heads, based on the intuition that the attended symbols in the input image should be consistent for all attention heads at each decoding step.
@@ 2.2 Deep learning-based HMER methods
Model architectures: Over the past few years, various image-to-text tasks have benefited from the encoder-decoder architecture, such as image captioning (28), scene text recognition (29), and handwritten text recognition (30). Inspired by these advances, Deng et al. (31) first introduced the encoder-decoder framework into the offline mathematical expression recognition task. Afterward, a series of important HMER methods have been proposed (2; 3), achieving excellent results on CROHME benchmarks. They are equipped with a convolution neural network (CNN) encoder and a gated recurrent unit (GRU) (32) decoder, where (3) uses a multi-scale DenseNet (33) encoder to replace the VGG (34) architecture in (2).
@@
Attention mechanisms: Zhang et al. (2) first introduced the coverage mechanism to solve the under-parsing and over-parsing problems in HMER. Under-parsing means that some parts of the image are omitted from the decoding, while over-parsing implies that certain areas are unnecessarily parsed multiple times in the decoding process. The coverage mechanism uses historical alignment information to prompt the attention module to focus on unattended regions. With the excellent performance of the coverage mechanism, it has been widely used in subsequent RNN-based HMER methods (8; 9; 3; 15; 37; 38). However, the coverage mechanism can not be applied directly to the Transformer decoder due to its parallel computing. CoMER (6) made it possible to compute coverage attention in a Transformer decoder without harming its parallel decoding nature and achieved the leading ExpRate on three CROHME benchmarks. Moreover, Zhang et al. (3) proposed a multi-scale attention module that deals with recognizing symbols in different scales. Similarly, Bian et al. (8) proposed an attention aggregation module to integrate multi-scale coverage attention. As a core component of the encoder-decoder framework, attention mechanisms have attracted much research interest in the HMER community.
Training strategies: Various weakly supervised information have been introduced to improve the performance of HMER methods. For example, Truong et al. (39) introduced information about the absence or presence of symbols to the encoder. Li et al. (9) proposed a weakly supervised counting module and jointly optimized HMER and symbol counting tasks. Regarding bidirectional training, Bian et al. (8) proposed ABM that consists of two parallel inverse decoders, which were enhanced via mutual distillation. Besides, BTTR (5) applied a bidirectional training strategy to the Transformer decoder for bidirectional language modeling. For data augmentation, Li et al. (37) proposed a scale augmentation method for HMER, which scales input images randomly while maintaining their aspect ratio. PAL (38) was proposed to use printed mathematical expressions as extra data to help train the model via paired adversarial learning, which maps handwritten features into printed ones.
@@ 3 Proposed method
In this section, we first describe the overall architecture of our model and then introduce the proposed attention guidance mechanism, including how self-guidance and neighbor-guidance are applied to the Transformer decoder.
@@ 3.1 Overall architecture
An overview of our model for handwritten mathematical expression recognition is presented in Fig. 3. The model comprises a CNN encoder and a Transformer decoder with the proposed attention guidance mechanism.
Figure 3: Handwritten mathematical expression recognition with the proposed attention guidance mechanism. The model is based on CoMER (6), which consists of a CNN encoder and a bidirectionally trained Transformer decoder. The bidirectional decoder receives input and produces output in two directions simultaneously. Attention guidance is applied to the cross-attention modules of the decoder, where attention weights are refined via a guidance map.
@@ 3.1.2 Transformer decoder
Transformer decoder has been widely used in sequence generation tasks (43; 44; 6) due to its effectiveness at capturing complex dependencies within the input sequence and between the input and output sequences. The Transformer decoder architecture is characterized by a stack of identical layers comprising a self-attention module, a cross-attention module, and a feed-forward layer, as shown in Fig 3.
@@
In addition to the original cross-attention block, an attention refinement module (ARM) (6) is adopted to solve the over-parsing problem, as is the convention of previous HMER methods (8; 9; 3; 15; 37; 38). ARM uses past alignment information (i.e., the accumulation of past attention maps) to refine attention weights. Given the correlation map $\mathbf{E}_{t}$ and the attention map $\mathbf{A}_{t}$ at each decoding step $t$, ARM refines the correlation map using a convolution function $\phi(\cdot)$:
$\mathrm{ARM}(\mathbf{E}_{t},\mathbf{A}_{t})=\mathbf{E}_{t}-\phi(\mathbf{C}_{t})$ (10)
@@ 3.2 Attention guidance mechanism
where $\mathbf{\hat{E}}$ denotes the refined correlations, $f_{\mathrm{proj}}(\cdot)$ represents a linear projection, and $\odot$ denotes the element-wise multiplication. Given a set of attention maps $\mathbf{A}^{G}\in\mathbb{R}^{T\times L\times h}$ as the guidance, $\mathbf{G}$ can be obtained by averaging the multiple heads or the convolution function $\phi(\cdot)$ in Eq. 10:
@@ 3.2.1 Self-guidance
Considering that the image areas of symbols attended by multiple heads should be consistent, we propose to refine raw correlations by eliminating inconsistencies of different attention heads. As shown in Fig. 5, given raw correlations $\mathbf{E}_{1},\ldots,\mathbf{E}_{h}$ for $h$ heads, the attention guidance can be derived as the concatenation of attention maps computed in Eq. 7. Afterwards, the guidance map $\mathbf{G}$ is obtained by Eq. 13:
@@
where $\mathbf{A}_{1},\ldots,\mathbf{A}_{h}$ are attention weights computed from different perspectives, since each $\mathrm{Head}_{i}$ learns its own set of linear projection matrices $\mathbf{W}_{i}^{Q},\mathbf{W}_{i}^{K}$ for queries and keys. Usually, correct activations are consistent across multiple heads otherwise they might be noise. $\mathbf{G}$ is obtained by finding inconsistencies through a convolution function $\phi(\cdot)$. The self-guidance module then generates residual correlations to obtain the refined correlations $\mathbf{\hat{E}}$ in Eq. 12:
@@
In this way, self-guidance can be readily applied to multi-head attention blocks. Self-guidance facilitates feature aggregation of cross-attention modules since misleading visual features are explicitly suppressed. However, it does not exploit external information except attention maps in the current attention module, which can be compensated by the neighbor-guidance in Sec. 3.2.2.
@@ 3.2.2 Neighbor-guidance
where $\mathbf{\hat{A}}^{L}_{t-1}\in\mathbb{R}^{L\times h}$ is the attention guidance, and $\mathrm{mean}(\cdot)$ represents the average operation of the $h$ heads, which can be regarded as seeking consensus among the attention maps. Similar to the self-guidance module, residual correlations are then generated to obtain the refined correlations $\mathbf{\hat{E}}^{l}_{t}$:
@@
where $\alpha$ is a factor for adjusting the guided correlations, which is set to 1$\sim$5 empirically. Note that all the components of neighbor-guidance do not require training, making it convenient to apply to stacked decoders. With the help of neighbor-guidance, the attention weights of the middle layers are refined, thus improving the attention quality of both the current and final layers.

### 05 Mar 2024 v2

@@ Abstract
Handwritten mathematical expression recognition (HMER) is challenging in image-to-text tasks due to the complex layouts of mathematical expressions and suffers from problems including over-parsing and under-parsing. To solve these, previous HMER methods improve the attention mechanism by utilizing historical alignment information. However, this approach has limitations in addressing under-parsing since it cannot correct the erroneous attention on image areas that should be parsed at subsequent decoding steps. This faulty attention causes the attention module to incorporate future context into the current decoding step, thereby confusing the alignment process. To address this issue, we propose an attention guidance mechanism to explicitly suppress attention weights in irrelevant areas and enhance the appropriate ones, thereby inhibiting access to information outside the intended context. Depending on the type of attention guidance, we devise two complementary approaches to refine attention weights: self-guidance that coordinates attention of multiple heads and neighbor-guidance that integrates attention from adjacent time steps. Experiments show that our method outperforms existing state-of-the-art methods, achieving expression recognition rates of 60.75% / 61.81% / 63.30% on the CROHME 2014/ 2016/ 2019 datasets.
@@ 1 Introduction
Handwritten mathematical expression recognition (HMER) aims to translate mathematical expression images into corresponding  sequences. HMER is a fundamental OCR task in educational scenarios, which has led to many real-world applications, including automatic scoring and online education. However, HMER is challenging because mathematical expressions are laid out in two dimensions, different from natural language texts including English, Chinese, etc. An HMER algorithm is expected to analyze the structure of an expression while recognizing the symbols (1). Over the past few years, HMER methods have benefited from deep neural networks and have made significant progress (2; 3; 4; 5; 6; 7; 8; 9; 10; 11; 12). Most prevalent methods treat HMER as a sequence-to-sequence task and follow the attention-based encoder-decoder framework. In this framework, the encoder first extracts the visual features of the input image as a sequence, and then the decoder generates the target sequence by attending to the encoded feature sequence (2; 4; 5) at each decoding step. The attention mechanism attends to image regions at each decoding step based on the hidden state of the last step.
@@
Beyond the hidden state, many HMER methods (2; 6; 3; 8; 9) exploit historical attention weights to prompt the attention module to focus on unattended image areas, called the coverage mechanism. The coverage mechanism is proposed to solve the over-parsing and under-parsing problems. Over-parsing occurs when a particular symbol region is parsed multiple times unnecessarily, whereas under-parsing happens when a symbol region is overlooked and not parsed at all (2). The coverage mechanism successfully alleviates over-parsing but has limitations in addressing under-parsing. Fig. 1 illustrates a series of attended image regions from a state-of-the-art coverage-based system (6). As we can see, the superscripts of “$j^{2}$” and “$q^{2}$” are activated at the same time. As a result, features of another “2” that should be decoded in the future are incorporated in the current context, causing “$q^{2}$” to be omitted in the decoding process.
@@
We refer to the above observation as the context leakage phenomenon: if the attention module attends to image areas that should be parsed at subsequent decoding steps, context from the future will be incorporated into the attention result, causing symbols between the target region and the incorrectly attended areas to tend to be omitted. The coverage mechanism cannot solve this phenomenon since it is only aware of past alignment information. Different from the multi-activation phenomenon in (13; 14), the context leakage phenomenon emphasizes the information leakage caused by the early activation of image areas that should be parsed after the target region.
To address this issue, we propose an attention guidance mechanism which aims to explicitly suppress attention weights in irrelevant areas and enhance ones in the target region. This mechanism comprises two complementary approaches: self-guidance and neighbor-guidance.
@@
Self-guidance aims to refine correlation maps by eliminating inconsistencies of different attention heads. In multi-head attention, each head learns its own set of linear projection matrices for queries, keys, and values. This independence allows the model to capture a richer variety of information from different perspectives. Although different heads attend to image areas in different subspaces, the attended symbol in the input image should be consistent across multiple heads. However, due to symbols with similar visual features and context, certain heads may attend to different symbols. The incorrectly attended symbols are the cause of the context leakage phenomenon. Therefore, inconsistent attention should be eliminated. We achieve this through a guidance map indicating inconsistencies among attention heads and a linear projection to adjust the guidance weights.
While self-guidance is performed in each single decoding step, attention from adjacent time steps can also be involved in the attention guidance mechanism. For HMER methods with a stacked decoder (15; 5; 6; 10), the query of the attention module in each decoder layer is determined by the output of its previous layer. Fig. 2 illustrates how a stacked decoder generates the final attention regions. As we can see, only the output layer focuses on the symbol to be generated. The middle layers first calculate attention weights based on the previous symbol instead of directly aligning with the current one. In other words, each symbol is aligned via its previously decoded neighbor. Different HMER methods (5; 6) with different decoder layers ($L\geq 2$) follow this process consistently. Moreover, a similar phenomenon has also been found in image captioning systems (16; 17).
@@
Based on the above observation, we propose another attention guidance approach, neighbor-guidance. Neighbor-guidance aims to leverage the final attention weights of the previous decoding step to enhance appropriate attention for middle layers at the current step. For example, in Fig. 2, attention weights for generating “$x$” in layer $L$ can be reused to refine the attention weights of layer $L-2$ at decoding step 2, where the three “$x$” are attended simultaneously and the context leakage phenomenon occurs. With the help of neighbor-guidance, appropriate attention weights of middle layers are enhanced, thus improving the attention quality of both the current and final layers. Neighbor-guidance is implemented similarly to self-guidance but changes the attention guidance from self-generated attention to the attention of the last decoding step and can be readily employed without training. Moreover, different from self-guidance, neighbor-guidance can be applied only to middle layers and facilitates information propagation between adjacent time steps.
@@
    We propose the phenomenon of context leakage, which cannot be solved by the widely used coverage mechanism. To solve this issue, we propose a general attention guidance mechanism to explicitly suppress attention weights in irrelevant areas and enhance ones in the target region.
@@
    We devise self-guidance to refine the raw attention by seeking consensus among multiple attention heads, based on the intuition that the attended symbol in the input image should be consistent for all attention heads at each decoding step.
@@ 2.2 Deep learning-based HMER methods
Model architectures. Over the past few years, various image-to-text tasks have benefited from the encoder-decoder architecture, such as image captioning (28), scene text recognition (29), and handwritten text recognition (30). Inspired by these advances, Deng et al. (31) first introduced the encoder-decoder framework into the offline mathematical expression recognition task. Afterward, a series of important HMER methods have been proposed (2; 3), achieving excellent results on CROHME benchmarks. They are equipped with a convolution neural network (CNN) encoder and a gated recurrent unit (GRU) (32) decoder, where (3) uses a multi-scale DenseNet (33) encoder to replace the VGG (34) architecture in (2).
@@
Attention mechanisms. As a core component of the encoder-decoder framework, attention mechanisms have attracted much research interest in the HMER community. Zhang et al. (2) first introduced the coverage mechanism to solve the under-parsing and over-parsing problems in HMER. Under-parsing means that some parts of the image are omitted from the decoding, while over-parsing implies that certain areas are unnecessarily parsed multiple times in the decoding process. The coverage mechanism uses historical alignment information to prompt the attention module to focus on unattended regions. With the excellent performance of the coverage mechanism, it has been widely used in subsequent RNN-based HMER methods (8; 9; 3; 15; 37; 38). However, the coverage mechanism cannot be applied directly to the Transformer decoder due to its parallel computing. CoMER (6) made it possible to compute coverage attention in a Transformer decoder without harming its parallel decoding nature and achieved the leading ExpRate on three CROHME benchmarks. Moreover, Zhang et al. (3) proposed a multi-scale attention module that deals with recognizing symbols in different scales. Similarly, Bian et al. (8) proposed an attention aggregation module to integrate multi-scale coverage attention.
Training strategies. Various weakly supervised information have been introduced to improve the performance of HMER methods. For example, Truong et al. (39) introduced information about the absence or presence of symbols to the encoder. Li et al. (9) proposed a weakly supervised counting module and jointly optimized HMER and symbol counting tasks. Regarding bidirectional training, Bian et al. (8) proposed ABM that consists of two parallel inverse decoders, which were enhanced via mutual distillation. Besides, BTTR (5) applied a bidirectional training strategy to the Transformer decoder for bidirectional language modeling. For data augmentation, Li et al. (37) proposed a scale augmentation method for HMER, which scales input images randomly while maintaining their aspect ratio. PAL (38) was proposed to use printed mathematical expressions as extra data to help train the model via paired adversarial learning, which maps handwritten features into printed ones.
@@ 3 Proposed method
In this section, we first describe the overall architecture of our HMER method based on (6) and then introduce the proposed attention guidance mechanism, which is an extension of the cross-attention module.
@@ 3.1 Overall architecture
An overview of the HMER model is presented in Fig. 3. The model comprises a CNN encoder and a Transformer decoder with the proposed attention guidance mechanism.
Figure 3: Handwritten mathematical expression recognition with the proposed attention guidance mechanism. The model is based on CoMER (6), which consists of a CNN encoder and a bidirectionally trained Transformer decoder. The bidirectional decoder receives input and produces output in two directions simultaneously. Attention guidance is applied to the cross-attention modules of the decoder, where attention weights are refined via the guidance map.
@@ 3.1.2 Transformer decoder
Transformer decoder has been widely used in sequence generation tasks (43; 44; 6) due to its effectiveness in capturing complex dependencies within the input sequence and between the input and output sequences. The Transformer decoder architecture is characterized by a stack of identical layers comprising a self-attention module, a cross-attention module, and a feed-forward layer, as shown in Fig 3.
@@
In addition to the original cross-attention block, an attention refinement module (ARM) (6) is adopted to solve the over-parsing problem, as is the convention of previous HMER methods (8; 9; 3; 15; 37; 38). ARM uses the accumulation of past attention maps to refine attention weights. Given the correlation map $\mathbf{E}_{t}$ and the attention map $\mathbf{A}_{t}$ at each decoding step $t$, ARM refines the correlation map using a convolution function $\phi(\cdot)$:
$\mathrm{ARM}(\mathbf{E}_{t},\mathbf{C}_{t})=\mathbf{E}_{t}-\phi(\mathbf{C}_{t})$ (10)
@@ 3.2 Attention guidance mechanism
where $\mathbf{\hat{E}}$ denotes the refined correlations, $f_{\mathrm{proj}}(\cdot)$ represents a linear projection, and $\odot$ denotes the element-wise multiplication. Given a set of attention maps $\mathbf{A}^{G}\in\mathbb{R}^{T\times L\times h}$ as the guidance, $\mathbf{G}$ can be obtained by averaging attention heads or the convolution function $\phi(\cdot)$ in Eq. 10:
@@ 3.2.1 Self-guidance
Considering that the image regions of symbols attended by multiple heads should be consistent, we propose to refine raw correlations by eliminating inconsistencies of different attention heads. As shown in Fig. 5, given raw correlations $\mathbf{E}_{1},\ldots,\mathbf{E}_{h}$ for $h$ heads, the attention guidance can be derived as the concatenation of attention maps computed in Eq. 7. Afterwards, the guidance map $\mathbf{G}$ is obtained by Eq. 13:
@@
where $\mathbf{A}_{1},\ldots,\mathbf{A}_{h}$ are attention weights computed from different subspaces, since each $\mathrm{Head}_{i}$ learns its own set of linear projection matrices $\mathbf{W}_{i}^{Q},\mathbf{W}_{i}^{K}$ for queries and keys. Correct activations of symbols should be consistent across multiple heads otherwise they might be noise. $\mathbf{G}$ is obtained by finding inconsistencies through the convolution function $\phi(\cdot)$. The self-guidance module then generates residual correlations to obtain the refined correlations $\mathbf{\hat{E}}$ in Eq. 12:
@@
In this way, self-guidance can be readily applied to multi-head attention blocks. Self-guidance facilitates feature aggregation of cross-attention modules since misleading visual features are explicitly suppressed. However, the raw attention is self-refined and lacks external guidance, which can be improved by the neighbor-guidance approach.
@@ 3.2.2 Neighbor-guidance
where $\mathbf{\hat{A}}^{L}_{t-1}\in\mathbb{R}^{L\times h}$ is the attention guidance, and $\mathrm{mean}(\cdot)$ represents the average of the $h$ heads, which can be regarded as seeking consensus among multiple attention heads. Similar to the self-guidance module, residual correlations are then generated to obtain the refined correlations $\mathbf{\hat{E}}^{l}_{t}$:
@@
where $\alpha$ is a factor for adjusting the guided correlations, which is set to $\left[1,5\right]$ empirically. Note that all the components of neighbor-guidance do not require training, making it convenient to apply to stacked decoders. With the help of neighbor-guidance, the attention weights of the middle layers are refined, thus improving the attention quality of both the current and final layers.